a pipe 30 cm long is open at both ends. find the fundamental frequency. velocity of sound in air is 330 m/sec

Grade:12

1 Answers

Arun
6 years ago
We can find fundamental frequency of of an open ended ( hollow) pipe from the relation ; Length L ( meters ) = 1/2 wave length ( meters ) and Frequency f ( Hertz Hz , meters/ second ) = velocity of sound , v ( 330 m/second ) / 2 L , length of pipe given L = 30 cm = 0.30 m , since L = 1/2 wave length , v , so 0.30 m = 0.5 wave length or wave length = 0.60 meter , putting the value of wave length = 0.60 meters and velocity of sound 330 meter/ second in the equation , f = 330 / 2 x 0.30 = 550 Hz